# A Researcher’s Guide to Empirical Risk Minimization

### v1

@@ Abstract
This guide develops high-probability regret bounds for empirical risk minimization (ERM). The presentation is modular: we state broadly applicable guarantees under high-level conditions and give tools for verifying them for specific losses and function classes. We emphasize that many ERM rate derivations can be organized around a three-step recipe—a basic inequality, a uniform local concentration bound, and a fixed-point argument—which yields regret bounds in terms of a critical radius, defined via localized Rademacher complexity, under a mild Bernstein-type variance–risk condition. To make these bounds concrete, we upper bound the critical radius using local maximal inequalities and metric-entropy integrals, recovering familiar rates for VC-subgraph, Sobolev/Hölder, and bounded-variation classes.
We also review ERM with nuisance components—including weighted ERM and Neyman-orthogonal losses—as they arise in causal inference, missing data, and domain adaptation. Following Foster and Syrgkanis (2023), we highlight that these problems often admit regret-transfer bounds linking regret under an estimated loss to population regret under the target loss. These bounds typically decompose regret into (i) statistical error under the estimated (optimized) loss and (ii) approximation error due to nuisance estimation. Under sample splitting or cross-fitting, the first term can be controlled using standard fixed-loss ERM regret bounds, while the second term depends only on nuisance-estimation accuracy. We also treat the in-sample regime, where nuisances and the ERM are fit on the same data, deriving regret bounds and giving sufficient conditions for fast rates.
@@ 1 Introduction
Empirical risk minimization (ERM) is a central tool of modern statistics and machine learning. In a typical problem, we choose $\hat{f}_{n}$ by minimizing the empirical risk $P_{n}\ell(\cdot,f)$ over a function class $\mathcal{F}$, and seek guarantees on the regret (or excess risk)
$P\ell(\cdot,\hat{f}_{n})-P\ell(\cdot,f_{0}),$
@@
where $f_{0}\in\arg\min_{f\in\mathcal{F}}P\ell(\cdot,f)$ is a population risk minimizer. While this principle is simple, deriving sharp rates in new settings can be technical. Proofs typically rely on tools from empirical process theory, such as uniform local maximal inequalities and concentration bounds (Van Der Vaart and Wellner, 1996; Geer, 2000; Giné and Koltchinskii, 2006; Wainwright, 2019). To avoid re-deriving regret bounds for each loss and function class, a common goal is to express them in terms of standard complexity measures—such as covering numbers and entropy integrals—for which a well-developed calculus is available.
@@
This guide presents high-probability regret analysis for ERM. We take a modular approach, isolating proof patterns and complexity bounds that are broadly applicable. The main message is that many ERM rate derivations can be organized around a three-step recipe (Geer, 2000; Abbeel and Ng, 2004; Bartlett et al., 2005; Koltchinskii, 2011; Wainwright, 2019): (i) a deterministic basic inequality; (ii) a uniform local concentration bound for the empirical-process term; and (iii) a fixed-point argument that turns the local bound into a regret rate. Under a mild Bernstein-type variance condition on the loss, we give a general theorem that bounds ERM regret in terms of the critical radius—equivalently, the localized Rademacher complexity—of the loss-difference class $\mathcal{F}_{\ell}:=\{\ell(\cdot,f)-\ell(\cdot,f_{0}):f\in\mathcal{F}\}$. This critical-radius viewpoint separates the statistical task of controlling the local complexity of $\mathcal{F}_{\ell}$ from the algebraic task of solving the associated fixed-point inequality to obtain a regret rate.
@@
To carry out step (ii) and compute critical radii in practice, we develop maximal inequalities that upper bound the local complexity of $\mathcal{F}_{\ell}$, and hence its critical radius, in terms of metric-entropy integrals (Van Der Vaart and Wellner, 2011; Lei et al., 2016; Wainwright, 2019). These results reduce many rate calculations to bounding covering numbers for $\mathcal{F}_{\ell}$. Covering numbers are well understood for many classical function classes and behave stably under basic operations, such as forming star-shaped hulls or applying Lipschitz transformations, which streamlines these calculations. This route recovers familiar critical radii and regret rates for VC-subgraph classes, Sobolev/Hölder classes, and bounded-variation classes.
[…]
@@
These notes are not intended as a comprehensive survey of ERM, nor do they aim to provide the sharpest possible regret bounds. Instead, they collect proof patterns and complexity bounds that we have found useful for deriving regret rates across problems. Our goal is to sit between two complementary literatures: the generality of localized Rademacher complexity arguments (Bartlett and Mendelson, 2002; Bousquet et al., 2003; Bartlett et al., 2005; Wainwright, 2019) and the practical convenience of uniform-entropy and maximal-inequality arguments (Van Der Vaart and Wellner, 1996, 2011). Similarly, Section 5 is meant as a complement to the orthogonal statistical learning framework of Foster and Syrgkanis (2023). For introductory background on ERM, helpful lecture-note treatments include Bartlett (2013), Guntuboyina (2018), and Sen (2018), while book-length introductions include Shalev-Shwartz and Ben-David (2014), Mohri et al. (2018), Wainwright (2019), and Koltchinskii (2011). For empirical-process treatments of ERM and M-estimation, see Van Der Vaart and Wellner (1996) and Geer (2000).
@@
Section 2 introduces the ERM setup and the high-probability language used throughout, and recalls how regret bounds translate into error bounds under curvature. Section 3 presents a high-level blueprint and intuition for regret proofs. Section 4.1 develops a general regret theorem in terms of localized Rademacher complexity and critical radii, while Section 4.2 provides entropy-based bounds that make critical-radius calculations concrete for common classes. Finally, Section 5 extends the framework to nuisance-dependent ERM, including weighted ERM, orthogonal losses, and in-sample nuisance estimation.
@@ 2.1 Problem setup and notation
2.2 High-probability guarantees (PAC bounds)
@@
To state finite-sample guarantees, we will often use PAC-style (probably approximately correct) bounds (Vapnik, 1999; Geer, 2000; Shalev-Shwartz and Ben-David, 2014; Mohri et al., 2018; Wainwright, 2019). Concretely, a PAC bound specifies a function $\varepsilon(n,\eta)$ of the sample size $n$ and a failure probability $\eta\in(0,1)$ such that, for every $n\in\mathbb{N}$ and every $\eta\in(0,1)$,
@@
The constant $C_{\eta}$ is left unspecified and may depend poorly on $\eta$. Moreover, this bound need not hold until the sample size exceeds $N_{\eta}$, which may be arbitrarily large. In contrast, a PAC bound gives an explicit finite-sample guarantee by providing an $\eta$-dependent error function $\varepsilon(n,\eta)$ (and, if needed, an explicit $N_{\eta}$). For example, the high-dimensional linear regression bound in (1) yields the $O_{p}$ statement $R(\hat{f}_{n})-R(f_{0})=O_{p}\bigl(\sigma^{2}s\log p/n\bigr)$. It also shows how the guarantee degrades as $\eta$ decreases: $\eta$ enters only through $\log(1/\eta)$, which grows logarithmically as $\eta\downarrow 0$.
Another advantage of PAC-style bounds over $O_{p}$ bounds is that they facilitate simultaneous control of many events. For example, suppose we wish to control the errors of a collection of ERMs $\{\hat{f}_{n,k}:k\in K(n)\}$, where the index set $K(n)$ grows with $n$. An $O_{p}$ statement for each $k$ is inherently pointwise: knowing $R(\hat{f}_{n,k})-R(f_{0,k})=O_{p}(\varepsilon_{n,k})$ for every fixed $k$ does not, by itself, imply that all bounds hold simultaneously when $|K(n)|\to\infty$. In contrast, if we have PAC bounds
@@
$\Pr\left(\forall k\in[K(n)]:\ R(\hat{f}_{n,k})-R(f_{0,k})\leq\varepsilon_{k}\left(n,\frac{\eta}{|K(n)|}\right)\right)\geq 1-\eta.$
Thus, PAC bounds let us tune $\eta$ to control the probability of any failure across a growing family of estimators, typically at the cost of an additional $\log|K(n)|$ term.
In regression, a common approach is to partition the covariate space $\mathcal{X}$ into $K(n)$ disjoint bins and perform empirical risk minimization within each bin (Wasserman, 2006; Takezawa, 2005). For example, if $\mathcal{F}$ consists of constant predictors and we use the squared error loss, then $\hat{f}_{n,k}$ is the constant function equal to the empirical mean of the $\approx n/K(n)$ responses in bin $k$. In this case, the maximum squared error over $K(n)$ independent empirical means, each based on $n/K(n)$ observations, is typically of order $K(n)\log K(n)/n$, a fact that can be used to derive regret bounds for histogram regression (Wasserman, 2006). Similar needs for uniform control arise beyond classical regression, for instance in sequential or iterative regression procedures in reinforcement learning, such as fitted value iteration and fitted $Q$-iteration (Munos, 2005; Munos and Szepesvári, 2008; Antos et al., 2007; van der Laan et al., 2025a; van der Laan and Kallus, 2025a, b). In these settings, regret is incurred at each iteration, and one must control the cumulative regret with high probability; see, for example, Lemma 3 and Theorem 2 of van der Laan and Kallus (2025a).
@@ 2.3 From regret to $L^{2}(P)$ error: curvature and strong convexity
It is common to derive high-probability bounds on the regret $R(\hat{f}_{n})-R(f_{0})$. In many problems, the regret is locally quadratic in $f-f_{0}$, so $\{R(\hat{f}_{n})-R(f_{0})\}^{1/2}$ behaves like a notion of distance between $\hat{f}_{n}$ and $f_{0}$. Nonetheless, we are often more interested in the $L^{2}(P)$ estimation error $\|\hat{f}_{n}-f_{0}\|$, which is typically more interpretable and is often needed for downstream analyses. In this section, we show how to translate regret bounds into $L^{2}(P)$ error bounds.
A standard sufficient condition is a local quadratic-growth (curvature) inequality: there exists $\kappa>0$ such that, for all $f\in\mathcal{F}$,
@@
When this holds, any high-probability bound of the form $R(\hat{f}_{n})-R(f_{0})\lesssim\varepsilon$ immediately yields $\|\hat{f}_{n}-f_{0}\|\lesssim\kappa^{-1/2}\sqrt{\varepsilon}$ at the same confidence level. This condition is natural when $R$ is twice differentiable with positive curvature near $f_{0}$, since a second-order Taylor expansion around $f_{0}$ implies quadratic growth in $\|f-f_{0}\|$ locally.
We now formalize this intuition. A function class $\mathcal{F}$ is convex if, for any $f,g\in\mathcal{F}$ and any $t\in[0,1]$, the convex combination tf+(1t)gtf+(1-t)g also lies in $\mathcal{F}$. We say that the risk function $R:\mathcal{F}\to\mathbb{R}$ is strongly convex at $f_{0}$ if there exists a curvature constant $\kappa\in(0,\infty)$ such that, for all $f\in\mathcal{F}$,
@@
where $\dot{R}_{f_{0}}(h):=\left.\frac{d}{dt}R(f_{0}+th)\right|_{t=0}$ denotes the directional derivative of $R$ at $f_{0}$ in direction $h$ (Equation 14.42 of Wainwright, 2019).

### v2

@@ Abstract
This guide provides a reference for high-probability regret bounds in empirical risk minimization (ERM). The presentation is modular: we begin with intuition and general proof strategies, then state broadly applicable guarantees under high-level conditions and provide tools for verifying them for specific losses and function classes. We emphasize that many ERM rate derivations can be organized around a three-step recipe—a basic inequality, a uniform local concentration bound, and a fixed-point argument—which yields regret bounds in terms of a critical radius, defined via localized Rademacher complexity, under a mild Bernstein-type variance–risk condition (Bartlett et al., 2005; Wainwright, 2019). To make these bounds concrete, we upper bound the critical radius using local maximal inequalities and metric-entropy integrals, thereby recovering familiar rates for VC-subgraph, Sobolev/Hölder, and bounded-variation classes.
We also study ERM with nuisance components—including weighted ERM and Neyman-orthogonal losses—as they arise in causal inference, missing data, and domain adaptation. Following Foster and Syrgkanis (2023), we highlight that these problems often admit regret-transfer bounds linking regret under an estimated loss to population regret under the target loss. These bounds typically decompose the regret into (i) statistical error under the estimated loss and (ii) approximation error due to nuisance estimation. Under sample splitting or cross-fitting, the first term can be controlled using standard fixed-loss ERM regret bounds, while the second depends only on nuisance-estimation accuracy. As a novel contribution, we also treat the in-sample regime, in which the nuisances and the ERM are fit on the same data, deriving regret bounds and showing that fast oracle rates remain attainable under suitable smoothness and Donsker-type conditions.
@@ 1 Introduction
Empirical risk minimization (ERM) is a central tool in modern statistics and machine learning. In a typical problem, we choose $\hat{f}_{n}$ by minimizing the empirical risk
$P_{n}\ell(\cdot,f):=\frac{1}{n}\sum_{i=1}^{n}\ell(Z_{i},f)$
over a function class $\mathcal{F}$. The corresponding population risk is $R(f):=E_{P}[\ell(Z,f)],$ and a basic goal is to obtain guarantees on the regret (or excess risk)
$R(\hat{f}_{n})-R(f_{0}),\quad f_{0}\in\arg\min_{f\in\mathcal{F}}R(f)$
@@
where $f_{0}$ is a population risk minimizer. Although the ERM principle is simple, deriving sharp regret rates in new settings can be technically involved. Proofs typically rely on tools from empirical process theory, such as uniform local maximal inequalities and concentration bounds (Van Der Vaart and Wellner, 1996; Geer, 2000; Giné and Koltchinskii, 2006; Wainwright, 2019). To avoid re-deriving regret bounds separately for each loss and function class, a common goal is to express them in terms of standard complexity measures, such as covering numbers and entropy integrals, for which a well-developed calculus is available.
@@
This guide presents a high-probability regret analysis of ERM. We take a modular approach that isolates proof patterns and complexity bounds with broad applicability. The main message is that many ERM rate derivations can be organized around a three-step recipe (Geer, 2000; Abbeel and Ng, 2004; Bartlett et al., 2005; Koltchinskii, 2011; Wainwright, 2019): (i) a deterministic basic inequality; (ii) a uniform local concentration bound for the empirical-process term; and (iii) a fixed-point argument that converts the local bound into a regret rate. Under a mild Bernstein-type variance condition on the loss, we give a general theorem that bounds ERM regret in terms of the critical radius—equivalently, the localized Rademacher complexity—of the loss-difference class
$\mathcal{F}_{\ell}:=\{\ell(\cdot,f)-\ell(\cdot,f_{0}):f\in\mathcal{F}\}.$
This critical-radius viewpoint separates the statistical task of controlling the local complexity of $\mathcal{F}_{\ell}$ from the algebraic task of solving the associated fixed-point inequality to derive a regret rate.
@@
To carry out step (ii) and compute critical radii in practice, we develop maximal inequalities that upper bound the local complexity of $\mathcal{F}_{\ell}$, and hence its critical radius, in terms of metric-entropy integrals (Van Der Vaart and Wellner, 2011; Lei et al., 2016; Wainwright, 2019). These results reduce many rate calculations to bounding covering numbers for $\mathcal{F}_{\ell}$. Covering numbers are well understood for many classical function classes and behave stably under basic operations, such as forming star-shaped hulls or applying Lipschitz transformations, which streamlines these calculations. This approach recovers familiar critical radii and regret rates for VC-subgraph classes, Sobolev/Hölder classes, and bounded-variation classes.
A second focus of this guide is ERM with nuisance components. Many modern estimators minimize empirical risks of the form $P_{n}\ell_{\hat{g}}(\cdot,f)$, where the loss depends on a nuisance component $\hat{g}$ estimated from the data, for example through inverse-probability weighting or pseudo-outcome regression. Following Foster and Syrgkanis (2023), nuisance components often do not require a new regret analysis: one can apply a standard regret bound to the ERM under the estimated loss and then use a regret-transfer inequality to control the error induced by using $\hat{g}$ in place of $g$, namely, the discrepancy between the population risks (and minimizers) under $g$ and $\hat{g}$. Building on Foster and Syrgkanis (2023), we develop regret-transfer bounds for weighted ERM with estimated weights. Under sample splitting or cross-fitting, these bounds combine with fixed-loss regret results to yield high-probability guarantees for nuisance-dependent ERM. We also treat the in-sample setting, in which the nuisances are estimated on the same data used to compute $\hat{f}_{n}$, and which therefore requires additional uniform concentration arguments. For suitably smooth optimization classes, such as Hölder or Sobolev classes, we show that oracle rates are attainable under Donsker-type conditions on the nuisance class.
@@
These notes are not intended to be a comprehensive survey of ERM, nor do they aim to provide the sharpest possible regret bounds. Instead, they collect proof patterns and complexity bounds that I have found useful for deriving regret rates across a range of problems. This guide is intended as a technical reference on regret analysis for researchers with some familiarity with empirical process theory (Van Der Vaart and Wellner, 1996; Sen, 2022) and empirical risk minimization (Wainwright, 2019). More specifically, the goal is to bridge two complementary literatures: the generality of localized Rademacher complexity arguments (Bartlett and Mendelson, 2002; Bousquet et al., 2003; Bartlett et al., 2005; Wainwright, 2019) and the practical convenience of uniform-entropy and maximal-inequality arguments (Van Der Vaart and Wellner, 1996, 2011). Similarly, Section 5 is intended to complement the orthogonal statistical learning framework of Foster and Syrgkanis (2023). These are working notes, and comments and corrections are very welcome.
Useful resources.
For introductory background on ERM, helpful lecture-note treatments include Bartlett (2013), Guntuboyina (2018), and Sen (2022), while book-length introductions include Shalev-Shwartz and Ben-David (2014), Mohri et al. (2018), Wainwright (2019), and Koltchinskii (2011). For empirical-process treatments of ERM and M-estimation, see Van Der Vaart and Wellner (1996), Geer (2000), and Sen (2022). I found Sen (2022) and Wainwright (2019) especially helpful as complementary resources.
@@
Section 2 introduces the ERM setup and the high-probability language used throughout, and recalls how regret bounds translate into error bounds under curvature. Section 3 presents a high-level blueprint and intuition for regret proofs. Section 4.1 develops a general regret theorem in terms of localized Rademacher complexity and critical radii, while Section 4.2 provides entropy-based bounds that make critical-radius calculations concrete for common classes. Section 5 introduces ERM with nuisance components, including weighted ERM and orthogonal losses with sample splitting and in-sample nuisance estimation. The appendix collects several local concentration inequalities for empirical processes and empirical inner products that support the main results and may be useful more broadly in ERM analyses.
@@ 2.1 Problem setup and notation
2.2 High-probability guarantees (PAC-style bounds)
@@
To state finite-sample guarantees, we will often use probably approximately correct (PAC)–style bounds (Vapnik, 1999; Geer, 2000; Shalev-Shwartz and Ben-David, 2014; Mohri et al., 2018; Wainwright, 2019). Concretely, a PAC-style bound specifies a function $\varepsilon(n,\eta)$ of the sample size $n$ and a failure probability $\eta\in(0,1)$ such that, for every $n\in\mathbb{N}$ and every $\eta\in(0,1)$,
@@
The constant $C_{\eta}$ is left unspecified and may depend poorly on $\eta$. Moreover, this bound need not hold until the sample size exceeds $N_{\eta}$, which may be arbitrarily large. In contrast, a PAC-style bound gives an explicit finite-sample guarantee by providing an $\eta$-dependent error function $\varepsilon(n,\eta)$ (and, if needed, an explicit $N_{\eta}$). For example, the high-dimensional linear regression bound in (1) yields the $O_{p}$ statement $R(\hat{f}_{n})-R(f_{0})=O_{p}\bigl(\sigma^{2}s\log p/n\bigr)$. It also shows how the guarantee degrades as $\eta$ decreases: $\eta$ enters only through $\log(1/\eta)$, which grows logarithmically as $\eta\downarrow 0$. PAC-style bounds can also be used to prove almost-sure convergence. In particular, if
$\sum_{n=1}^{\infty}\Pr\!\left(R(\hat{f}_{n})-R(f_{0})>\varepsilon(n,\eta_{n})\right)<\infty$
for some sequence $\{\eta_{n}\}_{n\in\mathbb{N}}$, then the Borel–Cantelli lemma implies that $R(\hat{f}_{n})-R(f_{0})\leq\varepsilon(n,\eta_{n})$ eventually almost surely.
Another advantage of PAC-style bounds over $O_{p}$ bounds is that they facilitate simultaneous control of many events. For example, suppose we wish to control the errors of a collection of ERMs $\{\hat{f}_{n,k}:k\in\mathcal{K}_{n}\}$, where the size $K(n):=|\mathcal{K}_{n}|$ of the index set $\mathcal{K}_{n}$ grows with $n$. An $O_{p}$ statement for each $k$ is inherently pointwise: knowing $R(\hat{f}_{n,k})-R(f_{0,k})=O_{p}(\varepsilon_{n,k})$ for every fixed $k$ does not, by itself, imply that all bounds hold simultaneously when $K(n)\to\infty$. In contrast, if we have PAC-style bounds
@@
$\Pr\left(\forall k\in\mathcal{K}_{n}:\ R(\hat{f}_{n,k})-R(f_{0,k})\leq\varepsilon_{k}\left(n,\frac{\eta}{K(n)}\right)\right)\geq 1-\eta.$
Thus, PAC-style bounds let us tune $\eta$ to control the probability of any failure across a growing family of estimators, typically at the cost of an additional $\log K(n)$ term.
In regression, a common approach is to partition the covariate space $\mathcal{X}$ into $K(n)$ disjoint bins and perform empirical risk minimization within each bin (Wasserman, 2006; Takezawa, 2005). For example, if $\mathcal{F}$ consists of constant predictors and we use the squared-error loss, then $\hat{f}_{n,k}$ is the constant function equal to the empirical mean of the $\approx n/K(n)$ responses in bin $k$. In this case, the maximum squared error over $K(n)$ independent empirical means, each based on $n/K(n)$ observations, is typically of order $K(n)\log K(n)/n$, a fact that can be used to derive regret bounds for histogram regression (Wasserman, 2006). Similar needs for uniform control arise beyond classical regression, for instance in sequential or iterative regression procedures in reinforcement learning, such as fitted value iteration and fitted $Q$-iteration (Munos, 2005; Munos and Szepesvári, 2008; Antos et al., 2007; van der Laan et al., 2025a; van der Laan and Kallus, 2025a, b). In these settings, regret is incurred at each iteration, and one must control the cumulative regret with high probability; see, for example, Lemma 3 and Theorem 2 of van der Laan and Kallus (2025a).
Switching between different ways of writing high-probability bounds
We state our results in PAC-style form, whereas some related works, such as Bousquet (2002), Bartlett et al. (2005), and Wainwright (2019), express analogous results as exponential tail bounds. For example, PAC-style bounds often take the form
$R(\hat{f}_{n})-R(f_{0})\lesssim\varepsilon_{n}+\frac{\log(1/\eta)}{n}$
with probability at least $1-\eta$, where $\varepsilon_{n}$ is the leading rate term and $\log(1/\eta)/n$ is the deviation term. Setting $\Delta=\log(1/\eta)/n$ yields the equivalent formulation that, with probability at least $1-e^{-n\Delta}$, the regret exceeds $\varepsilon_{n}$ by at most $\Delta$:
$R(\hat{f}_{n})-R(f_{0})\lesssim\varepsilon_{n}+\Delta.$
Thus, the PAC-style form immediately implies an exponential tail bound. Alternatively, following Wainwright (2019), we may choose $\eta=e^{-tn\varepsilon_{n}}$ for some $t>0$. Then $\log(1/\eta)/n=t\varepsilon_{n}$, so the above display implies that, with probability at least $1-e^{-tn\varepsilon_{n}}$,
$R(\hat{f}_{n})-R(f_{0})\lesssim(1+t)\varepsilon_{n}.$
In particular, $n\varepsilon_{n}$ can be viewed as an effective sample size for learning $f_{0}$, since it governs the exponent in the tail probability. More generally, if $n\varepsilon_{n}=O(1)$, then for any fixed confidence level $1-\alpha\in(0,1)$, we can choose $t$ large enough that $1-e^{-tn\varepsilon_{n}}\geq 1-\alpha$, and hence $R(\hat{f}_{n})-R(f_{0})=O_{p}(\varepsilon_{n})$. Integrating the tail bound over $t$ yields the mean regret bound $\mathbb{E}\{R(\hat{f}_{n})-R(f_{0})\}=O(\varepsilon_{n}+1/n)$ (Wainwright, 2019).
Each formulation parametrizes the same high-probability bound in a different way and may be more convenient in different applications. The first uses the failure probability $\eta$, the second uses the additive deviation $\Delta$, and the third uses the multiplicative inflation parameter $t$. The first is often convenient when combining many events via a union bound. The second and third more directly quantify the size of the deviation, and the third makes it especially easy to read off $O_{p}$ convergence rates.
@@ 2.3 From regret to $L^{2}(P)$ error: curvature and strong convexity
It is common to derive high-probability bounds on the regret $R(\hat{f}_{n})-R(f_{0})$. In many problems, the regret is locally quadratic in $f-f_{0}$, so $\{R(\hat{f}_{n})-R(f_{0})\}^{1/2}$ behaves like a distance between $\hat{f}_{n}$ and $f_{0}$. Nevertheless, the quantity of primary interest is often the $L^{2}(P)$ estimation error $\|\hat{f}_{n}-f_{0}\|$, which is typically more interpretable and is often needed for downstream analyses. In this section, we show how regret bounds can be translated into $L^{2}(P)$ error bounds.
A standard sufficient condition is a local quadratic-growth, or curvature, inequality: there exists $\kappa>0$ such that, for all $f\in\mathcal{F}$,
@@
When this holds, any high-probability bound of the form $R(\hat{f}_{n})-R(f_{0})\lesssim\varepsilon$ immediately implies $\|\hat{f}_{n}-f_{0}\|\lesssim\kappa^{-1/2}\sqrt{\varepsilon}$ at the same confidence level. This condition is natural when $R$ is twice differentiable with positive curvature near $f_{0}$, since a second-order Taylor expansion around $f_{0}$ yields local quadratic growth in $\|f-f_{0}\|$.
We now formalize this intuition. A function class $\mathcal{F}$ is convex if, for any $f,g\in\mathcal{F}$ and any $t\in[0,1]$, the convex combination tf+(1t)gtf+(1-t)g also belongs to $\mathcal{F}$. We say that the risk function $R:\mathcal{F}\to\mathbb{R}$ is strongly convex at $f_{0}$ if there exists a curvature constant $\kappa\in(0,\infty)$ such that, for all $f\in\mathcal{F}$,
@@
where $\dot{R}_{f_{0}}(h):=\left.\frac{d}{dt}R(f_{0}+th)\right|_{t=0}$ denotes the directional derivative of $R$ at $f_{0}$ in direction $h$; see Equation 14.42 of Wainwright, 2019.